# FGAIF: Aligning Large Vision-Language Models with Fine-grained AI Feedback

### v1

@@ 5.3 On Ablation Study (RQ2)
Table 4 shows the ablation study results of our FGAIF on several hallucination benchmarks. From this table, we have the following observations. 1) w/o-RLAIF performs terribly compared with FGAIF. It confirms the necessity of using RLAIF for modality alignment and hallucination mitigation in LVLMs. 2) FGAIF consistently outperforms w/o-Obj, w/o-Att, and w/o-Rel, across different evaluation metrics. This is reasonable because each reward model can provide feedback for one kind of hallucination. 3) FGAIF surpasses w-Coarse, denoting that the fine-grained reward models are more essential to align modalities in LVLMs compared with the traditional coarse-grained uni reward model.
@@ 5.4 On Case Study (RQ3)
To get an intuitive understanding of the hallucination mitigation capability of our model, we show two testing results of our method and LLaVA in Figure 3. Looking into the generated responses of the first sample, we can learn that by incorporating our fine-grained artificial intelligence feedback, our FGAIF is able to generate the faithful description for the input visual image, while the baseline cannot (e.g., the baseline generates “A seagull looking out at a lighthouse” and “a boat on the water” mistakenly). This intuitively demonstrates the necessity of considering the fine-grained feedback in reinforcement learning. A similar result can be found in the second sample.
@@ Acknowledgments
We thank Changchang Sun (Illinois Institute of Technology) and Jialu Li (University of North Carolina, Chapel Hill) for providing editing suggestions on the paper draft. We thank Yushi Hu (University of Washington) and Zhiqing Sun (Carnegie Mellon University) for the discussion on the initial idea.
Appendix A Prompts
We provide the prompt of annotating the consistency between the image and atomic fact in Figure 4. We also provide the prompt of atomic fact generation in Figure 5. In this prompt, we asked ChatGPT to generate three types of atomic facts: object existence, object attribute, and object relation. To get better performance on atomic fact generation, we added some samples in this prompt. You can refer to these broken-down samples to understand atomic facts.
Appendix B Experimental Settings
CHAIR is a framework to quantify object hallucination in image captions. This method compares objects generated in captions against the ground truth objects within the images. CHAIR assesses hallucination on two levels: sentence-level and instance-level. The sentence-level score, referred to as CHAIR, quantifies the proportion of captions that contain hallucinated content, whereas the instance-level score, CHAIR, measures the frequency of hallucinated objects relative to the total number of objects mentioned by the model. Our evaluation involves a randomly selected subset of 1,000 images from the MSCOCO validation set, allowing for an analysis of our model’s performance in minimizing object existence hallucination.
FaithScore is another framework to assess the accuracy and relevance of response generated by LVLMs. This innovative approach focuses on evaluating the consistency of atomic facts within the response against the depicted scenes in the input images. Different from CHAIR, FaithScore can demonstrate the model’s hallucination performance in terms of object existence, attribute, and relation. Our evaluation involves a randomly selected subset of 1,000 images from the MSCOCO validation set, allowing for an analysis of our model’s performance in mitigating object existence, attribute, and relation hallucination. It also provides an instance-level score F-Score and sentence-level score F-Score.
Appendix C Detailed Results
@@
| MiniGPT-4 | 79.7 | 80.2 | 52.5 | 69.7 | 73.0 | 62.2 | 65.2 | 70.4 | 67.8 | 74.5 | 60.8 |
| mPLUG-Owl | 54.0 | 68.4 | 95.6 | 50.9 | 66.9 | 98.6 | 50.7 | 66.8 | 98.7 | 67.2 | 97.6 |
| InstructBLIP | 88.6 | 89.3 | 56.6 | 79.7 | 80.2 | 52.5 | 65.2 | 70.4 | 67.8 | 80.0 | 59.0 |
| InstructBLIP | 88.7 | 89.3 | 55.2 | 81.4 | 83.5 | 62.6 | 74.4 | 78.5 | 69.0 | 83.7 | 62.2 |
| LLaVA | 50.4 | 66.6 | 98.8 | 49.9 | 66.4 | 99.4 | 49.7 | 66.3 | 99.4 | 66.4 | 99.2 |
| LLaVA | 73.7 | 78.8 | 72.3 | 73.6 | 78.2 | 71.0 | 67.2 | 74.4 | 77.8 | 77.1 | 73.7 |
| LLaVA-RLHF | 84.8 | 83.3 | 39.6 | 83.3 | 81.8 | 41.8 | 80.7 | 79.5 | 44.0 | 81.5 | 41.8 |
| LLaVA-RLHF | 85.2 | 83.5 | 38.4 | 83.9 | 81.8 | 38.0 | 82.3 | 80.5 | 40.5 | 81.9 | 39.0 |
| FGAIF | 87.0 | 86.7 | 45.9 | 84.0 | 83.7 | 48.1 | 79.6 | 79.9 | 50.9 | 83.4 | 48.3 |

### v2

@@ 5.3 On Ablation Study (RQ2)
Table 4 shows the ablation study results of our FGAIF on several hallucination benchmarks. From this table, we have the following observations. 1) w/o-RLAIF performs terribly compared with FGAIF. It confirms the necessity of using RLAIF for modality alignment and hallucination mitigation in LVLMs. 2) FGAIF consistently outperforms w/o-Obj, w/o-Att, and w/o-Rel, across different evaluation metrics. This is reasonable because each reward model can provide feedback for one kind of hallucination. 3) FGAIF surpasses w-Coarse, denoting that the fine-grained reward models are more essential to align modalities in LVLMs compared with the traditional coarse-grained uni reward model. 4) w/o-Obj performs worse than w/o-Att and w/o-Rel. This indicates that the object existence hallucination is most important. The potential reason is that there are more atomic facts of object existence, compared to the other hallucinations. 5) w/o-Att show a similar performance with w/o-Rel, which show w/o-Att has a similar importance with w/o-Rel.
To further evaluate the robustness of our reward model across responses with varying sub-segment counts, we constructed an additional test dataset. Specifically, we observed that the number of sub-segments in the training set ranged from [4, 16], while the original test set (comprising 500 samples) covered a range of [6, 15]. To assess model performance on longer responses, we collected an additional 200 samples with sub-segment counts between [15, 20]. Our evaluation results indicate an accuracy of 80.4% on the original test set and 79.2% on the newly constructed dataset. The comparable performance across different response lengths suggests that the impact of sub-segment length on model accuracy is minimal, demonstrating the robustness of our model.
During training, a fixed set of hallucination labels is introduced. A key question is whether the method can generalize to new objects that were not present in the training set. To assess the sensitivity of our approach to different object types, we followed prior works (Jiang et al., 2024; Yan et al., 2024) and constructed a dedicated out-of-the-distribution test set based on the Foggy dataset (Cordts et al., 2016). Specifically, we sampled 200 images from the Foggy test set to evaluate the reward model’s performance in this setting. The results show an accuracy of 76%, which is lower than the original test set but remains within an acceptable range. This suggests that while the model experiences some performance degradation when encountering unseen object types, it still maintains reasonable robustness.
@@ 5.4 On Case Study (RQ3)
To get an intuitive understanding of the hallucination mitigation capability of our model, we show two testing results of our method and LLaVA13B in Figure 5. Looking into the generated responses of the first sample, we can learn that by incorporating our fine-grained artificial intelligence feedback, our FGAIF is able to generate the faithful description for the input visual image, while the baseline cannot (e.g., the baseline generates “A seagull looking out at a lighthouse” and “a boat on the water” mistakenly). This intuitively demonstrates the necessity of considering the fine-grained feedback in reinforcement learning. A similar result can be found in the second sample.
5.5 Discussion
To better understand the quality of our AI-based feedback, we manually evaluated the accuracy of the atomic fact labels produced by the model. Specifically, we sampled 90 images from the MSCOCO validation set, generated corresponding answers along with their atomic facts, and manually annotated whether each atomic fact was consistent with the image. This was treated as a binary classification task—hallucination vs. non-hallucination. Based on this evaluation, we found that the accuracy of the AI-based feedback labeling was 85.07%, indicating that while the system is reasonably reliable, there is still room for improvement. These errors highlight the challenge of hallucination detection and suggest that enhancing the accuracy of AI-based feedback remains a promising direction for future work. Improving this component could further boost the effectiveness of our training signal and ultimately improve overall model performance.
@@ Acknowledgments
We thank the anonymous reviewers for valuable and insightful feedback. We thank Changchang Sun (Illinois Institute of Technology) and Jialu Li (University of North Carolina, Chapel Hill) for providing editing suggestions on the paper draft. We thank Yushi Hu (University of Washington) and Zhiqing Sun (Carnegie Mellon University) for the discussion on the initial idea. This research is supported in part by the National Science Foundation CAREER Grant IIS-2340435 and an Amazon Research Award. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of the U.S. Government.
Appendix A Detailed Results
@@
MiniGPT-47B | 79.7 | 80.2 | 52.5 | 69.7 | 73.0 | 62.2 | 65.2 | 70.4 | 67.8 | 74.5 | 60.8 |
mPLUG-Owl7B | 54.0 | 68.4 | 95.6 | 50.9 | 66.9 | 98.6 | 50.7 | 66.8 | 98.7 | 67.2 | 97.6 |
InstructBLIP7B | 88.6 | 89.3 | 56.6 | 79.7 | 80.2 | 52.5 | 65.2 | 70.4 | 67.8 | 80.0 | 59.0 |
InstructBLIP13B | 88.7 | 89.3 | 55.2 | 81.4 | 83.5 | 62.6 | 74.4 | 78.5 | 69.0 | 83.7 | 62.2 |
LLaVA7B | 50.4 | 66.6 | 98.8 | 49.9 | 66.4 | 99.4 | 49.7 | 66.3 | 99.4 | 66.4 | 99.2 |
LLaVA13B | 73.7 | 78.8 | 72.3 | 73.6 | 78.2 | 71.0 | 67.2 | 74.4 | 77.8 | 77.1 | 73.7 |
LLaVA-RLHF7B | 84.8 | 83.3 | 39.6 | 83.3 | 81.8 | 41.8 | 80.7 | 79.5 | 44.0 | 81.5 | 41.8 |
LLaVA-RLHF13B | 85.2 | 83.5 | 38.4 | 83.9 | 81.8 | 38.0 | 82.3 | 80.5 | 40.5 | 81.9 | 39.0 |
FGAIF7B | 87.0 | 86.7 | 45.9 | 84.0 | 83.7 | 48.1 | 79.6 | 79.9 | 50.9 | 83.4 | 48.3 |
Appendix B Error Cases
We show error cases for ChatGPT and LLaVA in Figure 6. In the error case for ChatGPT, ChatGPT should consider the white car as an object attribution rather than object existence. In another case for LLaVA, LLaVA misjudged the consistency between atomic fact and the images.
Appendix C PPO with Fine-Grained Rewards
The algorithm 1 shows in detail how PPO updates the policy LM and the value model , with fine-grained reward models .
Input: Initial policy model ; initial value model ; reward models trained from AI feedback;
task prompts ; hyperparameters , ,
Output: Updated policy model .
Appendix D Experimental Setups
The version of our ChatPT is gpt-3.5-turbo-0125. We set the number of beams and temperature to 1 and 0 to avoid randomness. The size of the LLaVA 1.5 feedback model is 13B.
Appendix E Abliation on Hallucination Labeler
To evaluate the dependency of our method on the backbone used for hallucination labeling, we followed your advice and conducted an additional experiment using an LVLM model—InternVL 2.5 (26B)—to label hallucinations. Using the same manually annotated set introduced earlier, we found that InternVL 2.5 achieved an accuracy of 89.47%, which is higher than that of the previously used LLaVA-based hallucination labeler. - Moreover, we re-ran our full pipeline using these improved hallucination labels and observed that better AI-based feedback led to consistently better performance across multiple benchmarks, as shown in Table 7. These results demonstrate that our method generalizes well across different vision-language backbones, and that the overall performance benefits from using stronger hallucination detectors.
Model | CHAIR_I ↓ | CHAIR_S ↓ | FaithScore ↑ | FaithScore_S ↑ | POPE F1 ↑ |
LLaVA_7B | 9.1 | 22.0 | 89.3 | 71.1 | 66.4 |
FGAIF_7B | 3.9 | 6.2 | 91.2 | 74.7 | 83.4 |
FGAIF_7B_new | 3.3 | 5.9 | 91.5 | 75.8 | 84.3 |
Appendix F Details of Sentence Splitter
The sub-sentence split process involves two steps: First, we use a spaCy model 222[https://spacy.io/usage/models to segment the input text into coarse-grained sentences. Then, each sentence is further split into sub-sentences using a custom heuristic. In particular, for each sentence splited in the spacy Model, a new sub-sentence is initiated after punctuation marks such as commas, semicolons, exclamation marks, or question marks.
Appendix G Detailed Discuss
While AI-based feedback significantly reduces human annotation cost, its imperfect accuracy (e.g., 85.07% for LLaVA-13B) introduces potential noise that may affect downstream learning. To understand how such noise propagates through the pipeline, we analyze its impact on two stages: reward model training and reinforcement learning. First, the reward model is trained on AI-generated labels, where label noise may cause incorrect supervision. However, we observe that the model still learns meaningful hallucination annotation, as demonstrated by strong alignment metrics (e.g., 91.2 FaithScore with noisy feedback). This suggests that the reward model exhibits robustness to moderate label noise. Second, during RL training, feedback errors may lead to unstable updates or suboptimal convergence. Empirically, we observe smooth reward curves and no mode collapse, indicating that the reward signal—while imperfect—still provides effective guidance. Overall, while feedback noise is non-negligible, our framework demonstrates empirical stability and effectiveness even with moderately noisy reward signals. Further increasing feedback accuracy (e.g., using InternVL 2.5) yields consistent gains (see Table 7), reinforcing the importance.
To further assess the impact of feedback model quality on the overall performance of our framework, we compared two vision-language models used during the feedback collection stage: LLaVA-13B, achieving 85.07% feedback accuracy, and InternVL 2.5, achieving a higher accuracy of 89.47%. As shown in Table 7, the model trained with InternVL-collected feedback (FGAIF_7B_new) consistently outperforms the LLaVA-based variant (FGAIF_7B) across all evaluation metrics. Specifically, we observed improvements in FaithScore (91.2 → 91.5), FaithScore_S (74.7 → 75.8), and POPE F1 (83.4 → 84.3), along with a reduction in hallucination rates (CHAIR_I: 3.9 → 3.3; CHAIR_S: 6.2 → 5.9). These results highlight that higher-quality feedback leads to better reward modeling and improved fine-tuning outcomes. Importantly, the modular nature of our pipeline allows easy substitution of backbone models, enabling the framework to continuously benefit from advancements in vision-language systems. Additionally, we replaced gpt-3.5-turbo-0125 with gpt-4o-2024-08-06 in our method and found that the performance remained consistent, achieving an F1 score of 83.5 on the POPE dataset. To further understand the underlying reasons, we sampled 10 data points annotated by both gpt-3.5 and gpt-4o and found a 98% overlap in the generated atomic facts (i.e., the number of overlapping atomic facts divided by the total number of unique atomic facts generated by both models). This indicates that both models produce highly similar atomic facts, leading to comparable performance.
our three-step pipeline—comprising feedback collection, reward model training, and reinforcement learning (RL) fine-tuning—aligns with the standard PPO Reinforcement Learning from Human Feedback (RLHF) framework, which also consists of these three stages. However, our approach introduces a significant enhancement by employing AI-based feedback mechanisms in place of traditional human annotations. This substitution substantially reduces the time and resources typically required for feedback collection, thereby accelerating the overall development process. - In terms of computational costs and training time, our method consists of reward model training and the RL stage, which is same to the other RLHF method [1]. Although our method need to train three reward models and compute rewards with three models in the RL stage, these processes can be run in parallel. Therefore, our approach does not introduce large additional computational overhead beyond the standard framework.
Appendix H More Qualitative Analysis
Figure 7 presents two additional examples that highlight typical hallucination patterns observed in vision-language models. In the first example, both models correctly identify the presence of a couch, but LLaVA-13B incorrectly predicts the presence of a chair, which is not visible in the image. In the second example, both models correctly identify the spoon, but LLaVA13B hallucinates a bowl, which is absent. These cases reveal a common hallucination tendency: co-occurrence-induced hallucinations. We observe that LLaVA-13B often hallucinates objects that frequently co-occur with the queried item in training data—for instance, assuming a bowl exists when a spoon is detected. This suggests that the model may rely heavily on dataset biases or conditional priors rather than grounding its answers purely in the visual evidence. In contrast, our FGAIF-aligned model correctly suppresses such hallucinations by aligning model behavior with image-grounded atomic facts. Our method explicitly penalizes misalignments during fine-tuning, which improves the model’s ability to distinguish what is visually present versus what is statistically likely. This showcases FGAIF’s strength in mitigating prior-driven hallucinations and promoting faithful visual understanding.